Matrix A is factored in the form PDP-1 where

A = PDP^{^{-1}} = \begin{bmatrix} -2 & 0& -1\\ 0& 1 & 2\\ 1 & 0 & 0 \end{bmatrix} \begin{bmatrix} 3 &0 &0 \\ 0 & -2 &0 \\ 0 & 0 & -2 \end{bmatrix}\begin{bmatrix} 0 &0 &1 \\ 2&1 &4 \\ -1&0 &-2 \end{bmatrix}

1. what are the eigenvalues of A and what are the dimensions of the corresponding eigenspaces?
Answer #1

Given - F-2 0-rsoal rooj] A = PDD To 1 2 || 0 -2 0 2 1 4 O II 0 0 Doo-2 1-2 0-17) 360) +0 (2) +0 (1) 3 (0)+0(1)+06) 3(1)+014(2-1) [62-1)(3-1)-001) - 0 [0/09-62-1903)-10 (910) f2-456)= → (2-1) 66+2x–31+12) -0 -10 (6-0) = 0 → 62-) (12-2) – 6 Joero 2-)бо Го о 21 o o o solue the metrin Axf]u=0 Гоо Гvg 2] [3-7 Ооо oo • J -J С. on t Izlo IttilsĽ210 -2-3 O ) tg = 3 farind Egon vector uz 0-4 ] = F2 -3 0 -107 1 0 0 3-3 = 75 a lot og all 19001217 door y saule the native२. **09-03 eigin quae ac* (3 003)